Question 13 (3 points) A company has a sales forecast for the following five months as shown in the table. If they have a beginning inventory of 225 units, what amount should be produced under a level plan in order for them to have an ending inventory of zero units at the end of the five month period? Month Forecast 825 600 650 550 475 anua Febru March A) 595 units per month 580 units per month 575 units per month D) 585 units per month
Answer #1

Total sales forecast of five months = 825+600+650+550+475 = 3100 units

Net sales forecast = Total sales forecast - beginning inventory = 3100-225 = 2875

Level production quantity = 2875/5 = 575 units per month

ANSWER: C) 575 units per month